OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} &\text{Find the number of integer values } \hspace{.33em}\\~\\ &\dfrac{15x^2+2x+1}{x^2-2x-1}\hspace{.33em}\\~\\ &\text{doesn't satisfy } \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

Hi math, Does that mean the number of integer values, which can not be written in the form:\[\dfrac{15x^2+2x+1}{x^2-2x-1}\]?

OpenStudy (mathmath333):

i mean the number of integer values the expression \(\dfrac{15x^2+2x+1}{x^2-2x-1}\) not take

OpenStudy (anonymous):

aha that's right, so we just need to find the range of function

OpenStudy (anonymous):

\[y=\dfrac{15x^2+2x+1}{x^2-2x-1}\]find \(x\) in terms of \(y\), find the domain of inverse

OpenStudy (anonymous):

@mukushla exact !

OpenStudy (mathmath333):

what u mean by x interms of y

OpenStudy (anonymous):

solve the equation . and deal with y as a parameter

OpenStudy (ehsan18):

Separate x i.e. find x = ? and replace y with x afterwards.

ganeshie8 (ganeshie8):

Following mukushla's hint \[\large y=\dfrac{15x^2+2x+1}{x^2-2x-1}\] cross multiplying and rearranging gives \[\large (15-y)\color{red}{x^2}+(2+2y)\color{red}{x}+y+1 = 0\] This is a quadratic in \(\color{red}{x}\) you may use the discriminant to find the domain

OpenStudy (mathmath333):

ok, d=8(y+1)(y-7)

ganeshie8 (ganeshie8):

how?

OpenStudy (mathmath333):

from this -> (2+2y)^2-4(15-y)(y+1)

ganeshie8 (ganeshie8):

Ohk.. Notice that \(D \ge 0\) for the quadratic equation to make sense in real numbers

OpenStudy (mathmath333):

\(y>=7 ,y<=-1\)

ganeshie8 (ganeshie8):

those are the values the given rational expression ever takes

OpenStudy (mathmath333):

is infinite numbers answer

ganeshie8 (ganeshie8):

answer should be {0,1,2,3,4,5,6}

OpenStudy (ehsan18):

no but these numbers satisfy the equation

OpenStudy (mathmath333):

oh ok thnx

OpenStudy (ehsan18):

what on earth the question says does not satisfy??

ganeshie8 (ganeshie8):